OpenStudy (anonymous):

A 73kg man weighs himself at the north pole and at the equator.. How much difference in weight is there, in Newtons?

OpenStudy (agent0smith):

You need to know the distance from the center of the earth to the north pole, and center of the earth to the equator. Then use: \[\Large F = G \frac{ m M }{ r^2 }\]M is the earth's mass, m is the man's mass. r is the distance.

OpenStudy (anonymous):

ok let me try that, one minute please

OpenStudy (anonymous):

would i also use the different gravity readings as well ?

OpenStudy (agent0smith):

What different gravity readings? That formula should account for them.

OpenStudy (anonymous):

its a difference of .03% between the two according to the net. but regardless I calculated 72.55 newtons. i entered the answer and it was not right. can you help.

OpenStudy (agent0smith):

How much difference in weight is there, in Newtons? the difference was 0.03%? Then how'd you get 72.55 newtons? Show your work...

OpenStudy (anonymous):

i know it doesn't make sense. G=6.67x10^-11 correct?

OpenStudy (agent0smith):

Correct. What values did you use for the two distances?

OpenStudy (anonymous):

earth = 6751 km equator = 6378

OpenStudy (anonymous):

converted to meters i just added 3 zeros

OpenStudy (anonymous):

this is how i set it up. (6.67X10^-11)*73*(5.972x10^24)/(4.53x10^13)

OpenStudy (anonymous):

answer i got was 641.90

OpenStudy (agent0smith):

That looks off.

OpenStudy (agent0smith):

The Earth's equatorial radius a, or semi-major axis, is the distance from its center to the equator and equals 6,378.1370 kilometers . Polar radius The Earth's polar radius b, or semi-minor axis, is the distance from its center to the North and South Poles, and equals 6,356.7523 kilometers These numbers are hugely different from your two earth = 6751 km equator = 6378

OpenStudy (anonymous):

yea thus why my calculations were wrong. the answer ended up being 2.5 newtons. i just don't see how though

OpenStudy (agent0smith):

That's about 5 newtons... but it'll depend on the exact numbers you used for the radii. I imagine they gave the radii to you...

OpenStudy (agent0smith):

Plus it depends what mass of earth you used....

OpenStudy (anonymous):

due to axial rotation of earth there is a slight decrease in g=earth's gravity in equator,and g=maximum at pole.we can go by this equation.\[g'(observed)=g-R \omega^2\cos ^2\theta \] at equator,\[\theta\]=0',at pole \[\theta \]=90' so at equator a man weighs=mg'=m(g-w^2R)=73X(9.8-.00192)=715N but at pole.9.8X73N=715.4N. the difference will be .4 N. hope this helps.

OpenStudy (agent0smith):

^the difference due to the radius of earth is much greater, though.